# On Function-Correcting Codes in the Lee Metric

### v2

@@ 5 FCLMCs for the locally functions
When $\lambda=2$, we say the function is a locally binary Lee function. As the optimal redundancy can be obtained by analyzing those message words for which function values are different. The following lemma distinguishes the message words inside a function ball. The proof follows a similar approach to that of the Hamming distance [9]. Therefore, we omit the proof.
@@
Let $f:\mathbb{F}_{2}^{k}\to Im(f)$ be a locally $(\lambda,\rho)_{L}$-function. Then there exists a map $\tau_{f}^{L}:\mathbb{F}_{2}^{k}\to[\lambda]$ such that for all $u,v\in\mathbb{F}_{2}^{k}$, if $f(u)\neq f(v)$ and $d_{L}(u,v)\leq\rho$, then $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$
Using the above lemma, we construct an $(f,t)$-FCLMC for locally $(\lambda,2t)_{L}$-functions. This will give an upper bound on the optimal redundancy in terms of the existence of classical codes in the Lee metric with prescribed parameters depending on $\lambda$ and $t$.
@@
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a locally $(\lambda,2t)_{L}$-function. Then, an upper bound on the optimal redundancy of an $(f,t)$-FCLMC is given by
@@
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a locally $(2,2t)_{L}$-function. Then $r_{f}^{L}(k,t)=N_{L}(2,2t)$.
@@
The Proof follows using Theorem 5.4 and Theorem 4.8. ∎
The above corollary suggests that determining the value $N_{L}(2,2t)$ will give the optimal redundancy of $(f,t)$-FCLMC for locally $(2,2t)_{L}$-functions.
Proposition 5.6.
For positive integers $t$ and $m\geq 2$, we have
Proof.
Thus, $N_{L}(2,2t)\leq s$.
Therefore, $N_{L}(2,2t)\geq s$. This completes the proof. ∎
Remark 5.7.
@@ 6 FCLMCs for the Lee weight and Lee weight distribution functions
In this section, we explore locality and optimal redundancy of FCLMCs for the Lee weight and the Lee weight distribution functions.
@@
The optimal redundancy of an $(f,t)$-FCLMC for the Lee weight function $f(u)=w_{L}(u)$ satisfies $r_{f}^{L}(k,t)\leq N_{L}(4t+1,2t)$.
@@
By Theorem 6.1, the Lee weight function $w_{L}(u)$ is a locally $(4t+1,2t)_{L}$-function. The rest follows from Theorem 5.4. ∎
We know that the maximum possible Lee weight of an element in $\mathbb{Z}_{m}$ can be $\left\lfloor\frac{m}{2}\right\rfloor$. In fact, for any $u\in\mathbb{Z}_{m}^{k}$, the Lee weight of $u$, $w_{L}(u)\in\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor\}=Im(w_{L})$. It is easy to see that for any $i,j\in\{0,1,2,\dots,k\}=Im(w_{H})$, there exist $u_{i}=(1^{i},0^{k-i}),u_{j}=(1^{j},0^{k-j})\in\mathbb{Z}_{m}^{k}$ such that $w_{H}(u_{i})=i$, $w_{H}(u_{j})=j$ and $d_{H}(u_{i},u_{j})=|i-j|$. In the Lee metric, finding such elements is non-trivial. In the following lemma, we construct such elements for the Lee metric.
@@
For all $i\in\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor\}=Im(w_{L})$, there exist an $u_{i}\in\mathbb{Z}_{m}^{k}$ such that $w_{L}(u_{i})=i$ and $d_{L}(u_{i},u_{j})=|i-j|$.
@@
Note that
$\left(m-\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)-(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)=$ $m-i+(a-1)\left\lfloor\frac{m}{2}\right\rfloor-i+(a-1)\left\lfloor\frac{m}{2}\right\rfloor$
$=$ $m+2(a-1)\left\lfloor\frac{m}{2}\right\rfloor-2i$
$\geq$ $m+2(a-1)\left\lfloor\frac{m}{2}\right\rfloor-2a\left\lfloor\frac{m}{2}\right\rfloor$
$=$ $m-\left\lfloor\frac{m}{2}\right\rfloor\geq 0.$
That is,
$m-\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)\geq i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor>0.$
Therefore,
$w_{L}\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)=\min\left\{i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor,m-\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)\right\}=i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor.$
@@
Note that $(a-1)\left\lfloor\frac{m}{2}\right\rfloor)<i\leq a\left\lfloor\frac{m}{2}\right\rfloor$. This implies that $\left\lfloor\frac{m}{2}\right\rfloor\leq i-a\left\lfloor\frac{m}{2}\right\rfloor\leq 0\leq\left\lfloor\frac{m}{2}\right\rfloor$. Therefore, again by Eq. 2.1,
@@
Let $f(u)=w_{L}(u)$ and $\eta=k\left\lfloor\frac{m}{2}\right\rfloor$. For any $k>\left\lceil\frac{t+1}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil$,
$r_{f}^{L}(k,t)\geq\frac{10t^{3}+30t^{2}+20t}{3(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}.$
@@
Let $f:\mathbb{Z}_{4}^{k}\to Im(f)$ be the Lee weight function. Then by Corollary 6.6, for any $k>\left\lceil\frac{t+1}{2}\right\rceil$,
@@
In [6, Corollary 4.2], the authors proved the same bound for the homogeneous metric. As the homogeneous metric and the Lee metric are the same over $\mathbb{Z}_{4}$. However, Corollary 6.6 gives the bound over $\mathbb{Z}_{m}$ for any positive integer $m$ with respect to the Lee metric.
@@
$W_{L}^{T}(u)=\left\lfloor\frac{w_{L}(u)}{T}\right\rfloor.$
The following theorem gives a value of $\lambda$ for which the Lee weight distribution function $W_{L}^{T}$ is a locally $(\lambda,\rho)_{L}$-function.
@@
Let $\rho$ be a positive integer. Then the Lee weight distribution function $W_{L}^{T}$ is a locally $\left(\left\lfloor\frac{2\rho}{T}\right\rfloor+2,\rho\right)_{L}$-function.
@@
The optimal redundancy of an $(f,t)$-FCLMC for the Lee weight distribution function $f(u)=W_{L}^{T}(u)$ satisfies $r_{f}^{L}(k,t)\leq N_{L}\left(\left\lfloor\frac{4t}{T}\right\rfloor+2,2t\right)$.
@@
By Theorem 6.8, the Lee weight distribution function $W_{L}^{T}(u)$ is a locally $\left(\left\lfloor\frac{4t}{T}\right\rfloor+2,2t\right)_{L}$-function. The rest follows from Theorem 5.4. ∎

### v3

@@ 5 FCLMCs for the locally functions
When $\lambda=2$, we say the function is a locally binary Lee function. As the optimal redundancy can be obtained by analyzing those message words for which function values are different.
Let $P$ be a nonempty set and let $\prec$ be a binary relation on $P$. We say that $\prec$ defines a total ordering on $P$ if the following conditions hold:
  1. 1.
    for any $\in P, aba\prec b$ and $b\prec a$ imply $a=b$;
  2. 2.
    whenever $a,b,c\in P$ satisfy $a\prec b$ and $b\prec c$, it follows that $a\prec c$;
  3. 3.
    for all $a,b\in P$, the elements $a$ and $b$ are comparable, that is, either $a\prec b$, $b\prec a$, or $a=b$.
A pair $(P,\prec)$ satisfying these properties is called a totally ordered set. A subset $I\subseteq P$ is called a contiguous block if, for any two elements $a,b\in I$ with $a\prec b$, any $c\in P$ with $a\prec c\prec b$ is also contained in $I$. The following lemma distinguishes the message words inside a function ball. The proof follows a similar approach to that of the Hamming distance [12]. Here, we provide the proof for completeness.
@@
Let $f:\mathbb{F}_{2}^{k}\to Im(f)$ be a locally $(\lambda,\rho)_{L}$-function. Suppose there is a total order on the image set $\text{Im}(f)$ such that $B_{L}^{f}(u,\rho)$ form a contiguous block for all $u\in\mathbb{Z}_{m}^{k}$. Then there exists a map $\tau_{f}^{L}:\mathbb{F}_{2}^{k}\to[\lambda]$ such that for all $u,v\in\mathbb{F}_{2}^{k}$, if $f(u)\neq f(v)$ and $d_{L}(u,v)\leq\rho$, then $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$.
Proof.
Let $(\text{Im}(f),\prec)$ be a total order set and $B_{L}^{f}(u,\rho)$ form a contiguous block for all $u\in\mathbb{Z}_{m}^{k}$. Denote $\text{Im}(f)=\{e_{1}\prec e_{2}\prec\dots,\prec e_{\eta}\}$ be the image set of $f$ ordered by $\prec$. Define a map $\gamma:\text{Im}(f)\to[\lambda]$ as $\gamma(e_{j})=1+(j\pmod{\lambda})$. Observe that for any contiguous block $I\subseteq\text{Im}(f)$ with $|I|\leq\lambda$, $\gamma$ is one-one on $I$. Define map $\tau_{f}^{L}:\mathbb{Z}_{m}^{k}\to[\lambda]$, $\tau_{f}^{L}(u)=\gamma(f(u))$. For $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$ and $d_{L}(u,v)\leq\rho$, that is, $f(u),f(v)\in B_{L}^{f}(u,\rho)$. Since $|B_{L}^{f}(u,\rho)|\leq\lambda$, $\gamma$ is one-one on contiguous block $|B_{L}^{f}(u,\rho)|$. Thus $\tau_{f}^{L}(u)=\gamma(f(u))\neq\gamma(f(v))=\tau_{f}^{L}(v)$. ∎
In the following proposition, we show that the contiguous block condition in Lemma 5.3 is always satisfied for the Lee weight function over $\mathbb{Z}_{m}^{k}$. Similarly, we can prove for the Lee weight distribution function as well.
Proposition 5.4.
Let $f(u)=w_{L}(u)$ be the Lee weight function over $\mathbb{Z}_{m}^{k}$ and $\rho$ be a positive integer. Then $B_{L}^{f}(u,\rho)$ forms a contiguous block for all $u\in\mathbb{Z}_{m}^{k}$.
Proof.
Let $f(u)=w_{L}(u)$ be Lee weight function over $\mathbb{Z}_{m}^{k}$. As $\text{Im}(f)\subset\mathbb{Z}$, let $\leq$ be the natural order on $\text{Im}(f)$ induced from $\mathbb{Z}$. Define a graph $G$ on $\mathbb{Z}_{m}^{k}$, $u,v\in\mathbb{Z}_{m}^{k}$ is connected if and only if $d_{L}(u,v)=1$. Trivially, for every $u\in\mathbb{Z}_{m}^{k}$, there exists $v\in\mathbb{Z}_{m}$ such that $d_{L}(u,v)=1$. Thus $G$ is connected.
For $u\in\mathbb{Z}_{m}^{k}$, let $B_{L}^{f}(u,\rho)\subseteq\text{Im}(f)$ be the Lee weight function ball of radius $\rho$ and centered at $u$. For $a,b\in B_{L}^{f}(u,\rho)$, let $w_{3}\in\text{Im}(f)$ with $a\leq c\leq b$. We claim that $c\in B_{L}^{f}(u,\rho)$, that is, $B_{L}^{f}(u,\rho)$ forms a contiguous block. There exist $u_{1},u_{2}\in\mathbb{Z}_{m}^{k}$ such that $w_{L}(u_{1})=a$ and $w_{L}(u_{2})=b$. Taking the shortest paths on $G$ from $u_{1}$ to $u$ and $u$ to $u_{2}$, we get
$u_{1}=x_{0},x_{1},\dots,x_{d_{1}}=u=y_{0},y_{1},\dots,y_{d_{2}}=u_{2}$
with $d_{L}(x_{i},x_{i+1})=1$ and $d_{L}(y_{j},y_{j+1})=1$. By triangle inequality, $d_{L}(u,x_{i})\leq i\leq d_{L}(u,u_{1})\leq\rho$, and $d_{L}(u,y_{j})\leq j\leq d_{L}(u,u_{2})\leq\rho$, that is $x_{i},y_{j}\in B_{L}(u,\rho)$ for all $0\leq i\leq d_{1}$ and $0\leq j\leq d_{2}$. Rewrite the sequence as $u_{1}=v_{0},v_{1},\dots,v_{n}=u_{2}$. Then, we have $|w_{L}(v_{i})-w_{L}(v_{i+1})|\leq 1$, implies that $a=w_{L}(v_{0}),w_{L}(v_{1}),\dots,w_{L}(v_{n})=b$ changes by at most one at each step. This implies that every integer between $a$ and $b$ will be attained. To see, let there is a $c$ with $a<c<b$ and there is no $v_{j}$ such that $w_{L}(v_{j})=c$. There are some $v_{i}$ and $v_{j}$ such that $w_{L}(v_{i})<c$ and $w_{L}(v_{j})>c$. Let $i^{\prime}=\max\{i|\ w_{L}(v_{i})<c\}$. Then $w_{L}(v_{i^{\prime}+1})>c$, and consequently, $|w_{L}(v_{i}^{\prime})-w_{L}(v_{i^{\prime}+1})|\geq 2$, which a contradiction. Thus, there is some $v_{i}$ such that $f(v_{i})=c$ whenever $a\leq c\leq b$. This completes the proof. ∎
Using Lemma 5.3, we construct an $(f,t)$-FCLMC for locally $(\lambda,2t)_{L}$-functions. This will give an upper bound on the optimal redundancy in terms of the existence of classical codes in the Lee metric with prescribed parameters depending on $\lambda$ and $t$.
@@
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a locally $(\lambda,2t)_{L}$-function satisfying the contiguous block condition in Lemma 5.3. Then, an upper bound on the optimal redundancy of an $(f,t)$-FCLMC is given by
@@
Let $f:\mathbb{Z}_{m}^{k}\to Im(f)$ be a locally $(2,2t)_{L}$-function. Then
$r_{f}^{L}(k,t)=\left\lceil\frac{2t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil.$
@@
Using Theorem 5.5, Theorem 4.7, and Proposition 4.9, we get the desired result. ∎
@@ 6 FCLMCs for the Lee weight and Lee weight distribution functions
In this section, we explore the locality and optimal redundancy of FCLMCs for the Lee weight and Lee weight distribution functions. Note that the Lee weight function and Lee weight distribution function satisfy the contiguous block condition in Lemma 5.3 with respect to the order induced from the set of integers $\mathbb{Z}$. Therefore, we can apply Theorem 5.5 to establish bounds on optimal redundancy of FCLMCs for the Lee weight and Lee weight distribution functions. In addition, Proposition 4.10 gives a construction of FCLMCs for these functions, attaining the corresponding bounds.
@@
The optimal redundancy of an $(f,t)$-FCLMC for the Lee weight function $f(u)=w_{L}(u)$ satisfies
$r_{f}^{L}(k,t)\leq(4t+1)\left\lceil\frac{t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil.$
@@
By Theorem 6.1, the Lee weight function $w_{L}(u)$ is a locally $(4t+1,2t)_{L}$-function. The rest follows from Theorem 5.5 and Proposition 4.10. ∎
We know that the maximum possible Lee weight of an element in $\mathbb{Z}_{m}$ can be $\left\lfloor\frac{m}{2}\right\rfloor$. In fact, for any $u\in\mathbb{Z}_{m}^{k}$, the Lee weight of $u$, $w_{L}(u)\in\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor\}=Im(w_{L})$. It is easy to see that for any $i,j\in\{0,1,2,\dots,k\}=Im(w_{H})$, where $w_{H}$ is the Hamming weight, there exist $u_{i}=(1^{i},0^{k-i}),u_{j}=(1^{j},0^{k-j})\in\mathbb{Z}_{m}^{k}$ such that $w_{H}(u_{i})=i$, $w_{H}(u_{j})=j$ and $d_{H}(u_{i},u_{j})=|i-j|$. In the Lee metric, finding such elements is non-trivial. In the following lemma, we construct such elements for the Lee metric.
@@
For all $i,j\in\{0,1,\dots,k\left\lfloor\frac{m}{2}\right\rfloor\}=Im(w_{L})$, there exist $u_{i},u_{j}\in\mathbb{Z}_{m}^{k}$ such that $w_{L}(u_{i})=i$, $w_{L}(u_{j})=j$ and $d_{L}(u_{i},u_{j})=|i-j|$.
@@
Since $(a-1)\left\lfloor\frac{m}{2}\right\rfloor<i\leq a\left\lfloor\frac{m}{2}\right\rfloor$, by Eq. 2.1, we have
$w_{L}\left(i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor\right)=i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor.$
@@
Note that $(a-1)\left\lfloor\frac{m}{2}\right\rfloor<i\leq a\left\lfloor\frac{m}{2}\right\rfloor$. This implies that $\left\lfloor\frac{m}{2}\right\rfloor\leq i-a\left\lfloor\frac{m}{2}\right\rfloor\leq 0\leq\left\lfloor\frac{m}{2}\right\rfloor$. Therefore, again by Eq. 2.1,
@@
Let $f(u)=w_{L}(u)$ and $\eta=k\left\lfloor\frac{m}{2}\right\rfloor$. For any $k\geq\left\lceil\frac{t+1}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil$,
$r_{f}^{L}(k,t)\geq\left\lceil\frac{10t^{3}+30t^{2}+20t}{3(t+2)^{2}\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil.$
@@
Let $f:\mathbb{Z}_{4}^{k}\to Im(f)$ be the Lee weight function. Then by Corollary 6.6, for any $k\geq\left\lceil\frac{t+1}{2}\right\rceil$,
@@
In [7, Corollary 4.2], the authors proved the same bound for the homogeneous metric. As the homogeneous metric and the Lee metric are the same over $\mathbb{Z}_{4}$. However, Corollary 6.6 gives the bound over $\mathbb{Z}_{m}$ for any positive integer $m$ with respect to the Lee metric.
For $a,b\in\mathbb{Z}$, define shifted modulo [6] as follows
$a\text{ smod }b\equiv(a-1)\pmod{b}+1\in\{1,2,\dots,b\}.$
For example, 4 $4\text{ smod }4\equiv$ 4 and $5\text{ smod }4\equiv 1$.
Now, we give a construction of FCLMCs for the Lee weight function over $\mathbb{Z}_{m}^{k}$. that attains the optimal redundancy in Corollary 6.6 for several cases.
Construction of FCLMC for the Lee weight function
For $t\geq 1$, let $p_{1},p_{2},\dots,p_{2t+1}$ be a code with the minimum Lee distance $2t+1$, that is, $d_{L}(p_{i},p_{j})\geq 2t$, and $p_{i}=p_{i\text{ smod }(2t+1)}$ for $i\geq 2t+2$. Define an encoding $\mathcal{E}$ over $\mathbb{Z}_{m}^{k}$ given by $\mathcal{E}(u)=(u,p_{w_{L}(u)+1})$. For $u,v\in\mathbb{Z}_{m}^{k}$ with $w_{L}(u)\neq w_{L}(v)$, let $w_{L}(u)=w_{1}\text{ smod }(2t+1)$, $w_{L}(v)=w_{2}\text{ smod }(2t+1)$.
Case 1 If $|w_{L}(u)-w_{L}(v)|\geq 2t+1$, then
$d_{L}(\mathcal{E}(u),\mathcal{E}(v))\geq d_{L}(u,v)=w_{L}(u-v)\geq|w_{L}(u)-w_{L}(v)|\geq 2t+1.$
Case 2 If $|w_{L}(u)-w_{L}(v)|<2t+1$, then $w_{1}\neq w_{2}$, that is, $p_{w_{L}(u)+1}\neq p_{w_{L}(v)+1}$. Thus, we have
$d_{L}(\mathcal{E}(u),\mathcal{E}(v))=d_{L}(u,v)+d_{L}(p_{w_{L}(u)+1},p_{w_{L}(v)+1})\geq 2t+1.$
Hence, by Definition 3.1, encoding $\mathcal{E}$ defines an $(f,t)$-FCLMC for the Lee weight function over $\mathbb{Z}_{m}^{k}$. In the following example, we construct $p_{i}$’s for various values of $t$.
Example 6.8.
(1) For $m\geq 4$ and $t=1$, let $p_{1}=(0,0)$, $p_{2}=(1,1)$, $p_{3}=\left(\left\lfloor\frac{m}{2}\right\rfloor,0\right)$, and $p_{i}=p_{i\text{ smod }3}$ for $i\geq 4$. Moreover, when $m=4$, the bound in Corollary 6.6 is tight.
(2) For $m\geq 5$ and $t=2$, let $p_{1}=(0,0,0),p_{2}=(2,2,0),p_{3}=(2,0,2),p_{4}=(0,2,2),p_{5}=(3,3,3)$, and $p_{i}=p_{i\text{ smod }5}$ for $i\geq 6$. Moreover, when $m=5$, the bound in Corollary 6.6 is tight.
@@
$W_{L}^{T}(u)=\left\lfloor\frac{w_{L}(u)}{T}\right\rfloor,u\in\mathbb{Z}_{m}^{k}.$
The following theorem gives a value of $\lambda$ for which the Lee weight distribution function $W_{L}^{T}$ is a locally $(\lambda,\rho)_{L}$-function for an arbitrary positive integer $\rho$.
@@
Let $\rho$ be a positive integer. Then the Lee weight distribution function $W_{L}^{T}$ over $\mathbb{Z}_{m}^{k}$ is a locally $\left(\left\lfloor\frac{2\rho}{T}\right\rfloor+2,\rho\right)_{L}$-function.
@@
The optimal redundancy of an $(f,t)$-FCLMC for the Lee weight distribution function $f(u)=W_{L}^{T}(u)$ satisfies
$r_{f}^{L}(k,t)=\begin{cases}0&\text{ if }T>k\left\lfloor\frac{m}{2}\right\rfloor\\ \left\lceil\frac{2t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil&\text{ if }T=k\left\lfloor\frac{m}{2}\right\rfloor,\end{cases}$
[…]
$r_{f}^{L}(k,t)\leq\left(\left\lfloor\frac{4t}{T}\right\rfloor+2\right)\left\lceil\frac{t}{\left\lfloor\frac{m}{2}\right\rfloor}\right\rceil.$
Moreover, Proposition 4.10 provides a construction of FCLMC achieving the above bound.
@@
If $T>k\left\lfloor\frac{m}{2}\right\rfloor$, then $W_{L}^{T}(u)=0$ for all $u\in\mathbb{Z}_{m}^{k}$, that is, $W_{L}^{T}$ is a constant function. Hence, no redundancy is required. If $T=k\left\lfloor\frac{m}{2}\right\rfloor$, then $W_{L}^{T}(u)\in\{0,1\}$, that is, $W_{L}^{T}$ is $(2,2t)_{L}$-function. By Corollary 5.6, we get the desired result. For $T<k\left\lfloor\frac{m}{2}\right\rfloor$, by Theorem 6.9, the Lee weight distribution function $W_{L}^{T}(u)$ is a locally $\left(\left\lfloor\frac{4t}{T}\right\rfloor+2,2t\right)_{L}$-function. The rest follows from Theorem 5.5 and Proposition 4.10. ∎